OpenStudy (el_arrow):

test the series for convergence or divergence. Use comparison or limit comparison test.

OpenStudy (el_arrow):

\[\sum_{n=1}^{\infty} \frac{n }{ n(n+2) }\]

OpenStudy (el_arrow):

do i do partial fraction decomposition here?

OpenStudy (rational):

whats stopping you from doing the obvious.. cancelling n top and bottom ?

OpenStudy (el_arrow):

nothing

OpenStudy (rational):

just double check there are no typoes

OpenStudy (el_arrow):

\[\sum_{n=1}^{\infty} \frac{ 1 }{ n+2 }\]

OpenStudy (el_arrow):

\[an = \frac{ 1 }{ n+2 }.....bn \frac{ 1 }{ n}\]

OpenStudy (el_arrow):

i mean bn = 1/n

OpenStudy (rational):

looks good, keep going..

OpenStudy (el_arrow):

\[\frac{ 1 }{ n+2 } > \frac{ 1 }{ n }......diverges?\]

OpenStudy (rational):

that inequality doesn't look correct n+2 > n that means 1/(n+2) < 1/n

OpenStudy (el_arrow):

yeah and if we plug in for example n=3......(1/5)<(1/3) so it converges

OpenStudy (el_arrow):

right?

OpenStudy (rational):

nope, SUM 1/n doesn't converge.

OpenStudy (el_arrow):

it doesnt? why not?

OpenStudy (rational):

\(\sum\limits_{n=1}^{\infty} \dfrac{1}{n}\) is harmonic series

OpenStudy (el_arrow):

oh yes it is harmonic

OpenStudy (el_arrow):

so then we can use either directive or limit comparison test right?

OpenStudy (rational):

yes use limit comparison test may be

OpenStudy (rational):

\(a_n = \dfrac{1}{n+1}\) and \(b_n = \dfrac{1}{n}\) \[\lim\limits_{n\to\infty}\dfrac{a_n}{b_n} = ?\]

OpenStudy (el_arrow):

sorry i lost connection

OpenStudy (el_arrow):

\[\lim_{n \rightarrow \infty} \frac{ \frac{ 1 }{ n+2 } }{ \frac{ 1 }{ n } } = \frac{ n }{ n+2 }\]

OpenStudy (rational):

yes what does that evaluate to

OpenStudy (el_arrow):

\[\frac{ \infty }{ \infty }\]

OpenStudy (rational):

use lhopital rule if u have to

OpenStudy (el_arrow):

then it will be \[\lim_{n \rightarrow \infty} \frac{ \ln n }{ n+2 } \]

OpenStudy (rational):

from where are you getting logarithms ?

OpenStudy (ikram002p):

Div

OpenStudy (el_arrow):

\[\frac{ \frac{ n }{ n } }{ \frac{ n }{ n }}+\frac{ \frac{ n }{ n } }{ \frac{ 2 }{ n } }\]

OpenStudy (el_arrow):

is that better?

OpenStudy (rational):

\[\lim_{n \rightarrow \infty} \frac{ \frac{ 1 }{ n+2 } }{ \frac{ 1 }{ n } } = \lim_{n \rightarrow \infty} \frac{ n }{ n+2 } = \lim_{n \rightarrow \infty} \dfrac{1}{1} = 1\]

OpenStudy (el_arrow):

so l'hopital's rule is just like taking the derivative?

OpenStudy (rational):

yes http://mathworld.wolfram.com/LHospitalsRule.html

OpenStudy (el_arrow):

i kindy forgot how to do it hehehe....

OpenStudy (el_arrow):

i got this other problem its the last one \[\sum_{n=1}^{\infty} \frac{ 1+\sqrt{n} }{ n}\]

OpenStudy (el_arrow):

just want to see if i do it right

OpenStudy (el_arrow):

\[an = \frac{ 1+\sqrt{n} }{ n }.....bn=\frac{ \sqrt{n} }{ n }=\frac{ n^\frac{ 1 }{ 2 } }{ n }=n^\frac{ 1 }{ 2 }\]

OpenStudy (rational):

use direct comparison test with harmonic series maybe..

OpenStudy (rational):

for \(n\gt 0 \) we have \[\dfrac{1+\sqrt{n}}{n} \gt \dfrac{1}{n}\] since the harmonic series \(\sum \dfrac{1}{n}\) diverges, the series \(\sum \dfrac{1+\sqrt{n}}{n}\) also diverges/

OpenStudy (el_arrow):

how do you know to use direct comparison test with harmonic series?

OpenStudy (rational):

do you see that \(\frac{1+\sqrt{n}}{n}\) is greater than \(\frac{1}{n}\) for all \(n\gt 0\) ?

OpenStudy (el_arrow):

oh yes so thats how you test it to see if its harmonic

OpenStudy (rational):

good, but we're not testing if it is harmonic or not we're just showing the terms in our series are GREATER than the terms in harmonic seris

OpenStudy (el_arrow):

so if its greater we use direct comparison test

OpenStudy (dan815):